OpenStudy (anonymous):

a 30-in tv set also has a width of 24 inches. what is the height

OpenStudy (cwrw238):

use pythagoras 30 in = hypotenuse of right angled triangle so 30^2 = 24^2 + x^2 where x = height of tv screen

OpenStudy (anonymous):

sorry dont quit understand

OpenStudy (anonymous):

would it be 18

OpenStudy (lilg132):

yes 18

OpenStudy (anonymous):

thanks

OpenStudy (lilg132):

30^2 = 900 24^2 = 576 900 = 576 + x^2 900 - 576 = 324^2 324^2 = x^2 square root of 324 = 18 x = 18

OpenStudy (anonymous):

thanks i did it a li that way but got the same answer